When the rate of blood flow in the aorta is 5 liter/min, the velocity
of the blood in the capillaries is about 0.33 mm/sec. If the average
diameter of a capillary is 0.008 mm, calculate the number of capillaries
in the circulatory system.
.
.
.
answer key : N = 7,5 x 10⁴
Answer #1

Answer:

Given

  • Rate of blood flow in the aorta is 5 liter/min,

  • Velocity of the blood in the capillaries is about 0.33 mm/sec.

  • Average diameter of a capillary is 0.008 mm,

  • Calculate the number of capillaries in the circulatory system.

We can use the below formula as

(Aorta) A1v1 = A2v2 (Capillaries)

Aorta A1v1 = 5 litre/ min = 5 *cubic meter / minute = 5*100 *100*100 mm / 60 sec = 83333.33 cubic mm / sec

V2 of capillaries = 0.33 mm/sec

A2 of capillaries = π r *r = 3.14 * 0.004 * 0.004 = 0.00005024 square mm

A2 *V2 = 0.33 * 0.00005024 = 0.000016579 cubic mm

Thus the total number of capalliaries = A1v1 / A2V2

= 83333.33 / 0.000016579

=5026438868 nos.

= 5.026 * E9